# Serving Our AI Overlords: Redesigning Data Systems for Agents

### v1

@@ 2. Case Studies
We use the BIRD text2SQL benchmark (Li et al, 2023) in our first study. We wanted to explore if present-day LLMs benefit from increasing the number of requests—in parallel or in sequence. We used DuckDB as our backend, and GPT-4o-mini and Qwen2.5-Coder-7B-Instruct as two LLMs. To first evaluate parallel requests, we simulated the behavior of an LLM agent “in charge,” with a number of “field” agents each independently attempting the task, followed by the agent-in-charge picking one among the corresponding solutions. We plot the average success rate versus the number of LLM attempts in Figure 1(a). To instead evaluate sequential requests, we had a single LLM agent issue queries until it was satisfied and once again plot the success rate versus the number of steps taken in Figure 1(b). We find that:
@@ 4.2. From Data Systems to Agents
In addition to simply answering probes, data systems should steer agents towards better probes, which in turn can lead to improved efficiencies. In this way, the data system acts in a more proactive (Zeighami et al, 2025) manner, akin to how a data engineer or administrator may assist data analysts in satisfying information needs as efficiently as possible. This information can be provided in addition to, or in lieu of the answers to the probes, in natural language. This steering can serve two purposes: (1) helping agents by providing auxiliary data-centric information the data system finds relevant, as a side-channel, and (2) providing feedback to agents on efficiency and costs to assist the agents in designing their probes. We envision sleeper agents within the data system that are invoked on-demand to gather information in parallel with answering probes, to be returned in addition to probe answers, as we discuss below.
Auxiliary Information. As we saw in Table 1, providing grounding hints or feedback can reduce the number of probes agents need to complete their tasks. We envision sleeper agents tasked with identifying and providing such hints as auxiliary information along with answers. For example, the sleeper agent could find and share other related tables—to be either joined with (as in join discovery, e.g., (Sarma et al, 2012)), or replacing the current table as the focus of analysis, especially if the current table proves irrelevant. Or rather than the agent having to guess why they got an empty result, the sleeper agent can provide feedback reminiscent of why-not provenance (Cheney et al, 2009), e.g., the probe assumed that states were encoded with two letter acronyms like “CA”, but instead they are listed out in entirety.
@@ 5.1. Supporting Exploration
Our agentic probes will consist of exploratory queries to establish grounding. Some explorations will inevitably be cast in natural language (NL) as agents lack knowledge about the underlying databases (e.g., “how to find out how many tables are stored?”) with others expressed using SQL (e.g., SELECT count(*) FROM
information_schema.tables
in PostgreSQL). Today’s databases are not designed to answer NL queries. The probe optimizer in our agent-first data system must therefore orchestrate the mix of NL and SQL queries by utilizing different agents at the scale of probes.
To illustrate, consider identifying the stores that show an increasing sales trend. Our agents will first need to find out which tables are used to store sales data. A straw person probe execution plan is to pose NL questions to a web search agent to discover how to look up table schema, and execute the found queries on our database. While these are simple queries on our database’s metadata tables, the outputs returned from such queries often contain lots of unnecessary information. For instance, PostgreSQL maintains hundreds of internal tables and indexes even without any user table defined. Coupled with the user tables, the results can easily grow to thousands of rows. Feeding all the rows to our query formulation agent is simply a waste of its limited context length.
Subsequently, to discover what constitutes an increasing trend, one strategy is to find examples of “trend queries” using NL with a web search agent, then feed the returned information to a query formulation agent to translate into SQL. We will likely get lots of example queries online, and our database will be bombarded with lots of inapplicable queries (e.g., they refer to non-existent tables, or identify the wrong trend). Worse yet, all such explorations will be mixed with other agents formulating solutions. With today’s data systems, we have no means to identify which queries are part of agentic exploration (and hence do not need to be evaluated completely). We envision that our probe optimizer will prioritize queries based on their phases (i.e., a form of admission control). Furthermore, we will store previously gleaned information using our agentic memory store to avoid repeated querying of the same information, and train agents to query our memory store instead of including such information as part of the prompt each time.
@@ 5.2.1. Intra-Probe Optimization
Deciding on What to Execute. The database must first decide what queries to run and to what degree of approximation, taking the probe and its briefs into account. This requires the database to reason about the data and probe semantics, including the agent’s goals and phases. To do so, the database can use semantic query and data understanding to check if they match user’s intents, and prune away queries it deems not semantically meaningful. For instance, during the exploration phase, the database can examine the projected columns in probe to see if they are relevant to the user’s intent, and if not prune such columns, or the probes away as a whole. Moreover, the database can compare probes within a batch, guided by probe’s briefs that may have specified approximation needs across probes. The database can then make cost estimates and compare information gain from the probes to decide which probes are more helpful and/or cheaper. For example, given two probes $P$ and $P^{\prime}$ the database can prune $P^{\prime}$ away if rows that would be returned by $P^{\prime}-P$ are deemed irrelevant to the agent’s goal. This is reminiscent of prior work on pruning queries as part of visualization recommendation (et al, 2015), and deciding query equivalence as part of query synthesis given user provided input/output examples (Zloof, 1975; Wang et al, 2017), although the scale of queries to compute the differences will be much larger in agentic workloads. Finally, the database can take the agent’s phase into account; for example, return coarse grain approximations during exploration, but more accurate answers during solution formulation. Beyond pruning queries, we envision agents will be able to examine other internal database states (e.g., buffer pool, outputs of query operators) to determine if it should continue with query evaluation, or move on to the next turn.
@@
Efficient Execution. As mentioned in Sec. 2, probes have substantial redundancy that we can exploit by sharing computation across them. Multi-query optimization (et al, 2012; Sellis, 1988), approximate query processing (Garofalakis et al, 2001) and caching of partial query results can be used to improve efficiency. However, there are new unique challenges. For example, different probes will have different approximation requirements and may be accompanied with termination criteria (a function that can be evaluated on partial results to know if they are sufficient, see Sec. 4), which makes it more difficult to reason about their semantics and what can be shared. Moreover, the database can incrementally evaluate queries, reminiscent of incremental query processing (Chandramouli et al, 2014), but with the new challenge of decision making across them; e.g., the database must decide which probe is the most useful to the agent and provide higher accuracy for that probe first before increasing accuracy for other probes. Finally, query planning and processing can be done jointly with optimization, e.g., the database can re-evaluate its decisions on what queries to run, or increase its level of approximation for some query during planning or processing as it obtains more information.
@@ 5.2.2. Inter-Probe Optimizations
Deciding on What to Execute. Besides the strategies discussed in Section 5.2.1, the database can consider all interactions with the agent to decide what queries to run. First, it can decide on queries to run based on whether they provide any new information given past queries answered. For example, when given probes $P$ and $P^{\prime}$ by the agents across consecutive turns, if the output between $P$ and $P^{\prime}$ is expected to not provide new information to the agent—e.g., $P^{\prime}$ adds new columns that are non-relevant to the agent’s goal—then $P^{\prime}$ can be dropped. Furthermore, the database can decide what queries to run in order to minimize the number of future follow-up probes. For example, based on the agent’s goal specified in the probe briefs, it can run a query it finds maximally useful to the agent exactly and to completion rather than approximately even if the current query may take longer, expecting that the extra computation upfront will reduce total runtime across future interactions with the agent.
@@ 6. Indexing, Storage, and Transactions
On the independence front, data systems treat queries as unrelated, such that concurrent access (specifically writes) from these queries must be isolated from each other. While this simplifies application logic and ensures consistency, these mechanisms prevent cooperative sharing of state with rare exceptions (Gupta et al, 2011). Instead of isolation, agentic workloads demand a more cooperative model—one that can safely share intermediate state across different probes, many of which are likely to be similar.
@@ 6.1. Agentic Memory Store
Updates to the Store. A separate concern is how this memory store is maintained during updates. Updates could be in the form of new probes being executed that may provide new information that augments or supersedes existing ones. Or, it could be to the underlying data or metadata, necessitating updates to any related information in the agentic memory. For example, if there is a schema update, the results of a prior probe that used that table may no longer be relevant. One approach is to allow this memory store be inconsistent with the data/metadata, and instead be updated by any new probes that discover that the information is stale. However, the downside is that the stale information may lead a new probe to make a mistake. For example, suppose the agentic memory indicates that the only relevant sales information can be found in three tables, but after that, additional relevant tables were added; here, new probes may end up returning incorrect results. Additional challenges emerge in supporting access control for multiple users. For example, agents acting on different users’ behalf could ask similar questions (e.g., ”Where is the employee’s availability stored?”). Sharing answers across such agents boosts efficiency—but raises privacy concerns, especially in the aggregate (Popa et al, 2011). Addressing these challenges will need to draw inspiration from work on knowledge bases as well as schema evolution.
@@ 6.2. Performing Branched Updates
Branch Isolation. Existing models of branching consistency, developed in the weak consistency era, e.g., in Bayou, Dynamo, or Tardis (Petersen et al, 1996; DeCandia et al, 2007; Crooks et al, 2016), can offer inspiration. However, agentic speculation goes further: multiple agents may create forks that must eventually reconcile—not just with the mainline, but with each other. This requires new models of multi-agent, multi-version isolation. Most branches will be similar—e.g., same schema, 90% identical data—but isolation requires that their effects remain logically separate.
Efficient forking and rollbacks. Naively duplicating entire databases per branch is prohibitively expensive and inefficient, making support for efficient forking crucial. Industrial systems like Neon (Neon, [n. d.]), Aurora (Verbitski et al, 2017) and academic systems like Tardis (Crooks et al, 2016) adopt copy-on-write approaches to lazily clone state. However, these are still far from what is needed for massive speculation. We need new concurrency mechanisms that exploit similarity across branches and preserve logical isolation (no cross-contamination), to enable massive parallel forking. This is analogous to MVCC on steroids: forking possibly thousands of near-identical snapshots and rolling back all but one. Unlike traditional data systems, where rollbacks are rare, we require ultra-fast rollbacks (i.e., fast aborts for failed branches).
@@ 7. Conclusion
We described our vision for a data system to support emerging agentic workloads. These workloads involve massive speculative probing, characterized by a high-throughput redundant mix of discovery and validation, specified by a combination of queries and natural language. We present an architecture for such redesigned data systems, and discuss emergent research challenges.

### v2

# Supporting Our AI Overlords: Redesigning Data Systems to be Agent-First

@@ 2. Case Studies
We use the BIRD text2SQL benchmark (bird) in our first study. We wanted to explore if present-day LLMs benefit from increasing the number of requests—in parallel or in sequence. We used DuckDB as our backend, and GPT-4o-mini and Qwen2.5-Coder-7B-Instruct as two LLMs. To first evaluate parallel requests, we simulated the behavior of an LLM agent “in charge,” with a number of “field” agents each independently attempting the task, followed by the agent-in-charge picking one among the corresponding solutions. We plot the average success rate versus the number of LLM attempts in Figure 1(a). To instead evaluate sequential requests, we had a single LLM agent issue queries until it was satisfied and once again plot the success rate versus the number of steps taken in Figure 1(b). We find that:
@@ 4.2. From Data Systems to Agents
In addition to simply answering probes, data systems should steer agents towards better probes, which in turn can lead to improved efficiencies. In this way, the data system acts in a more proactive (zeighami2025llm) manner, akin to how a data engineer or administrator may assist data analysts in satisfying information needs as efficiently as possible. This information can be provided in addition to, or in lieu of the answers to the probes, in natural language. This steering can serve two purposes: (1) helping agents by providing auxiliary data-centric information the data system finds relevant, as a side-channel, and (2) providing feedback to agents on efficiency and costs to assist the agents in designing their probes. We envision sleeper agents within the data system that are invoked on-demand to gather information in parallel with answering probes, to be returned in addition to probe answers, as we discuss below.
Auxiliary Information. As we saw in Table 1, providing grounding hints or feedback can reduce the number of probes agents need to complete their tasks. We envision sleeper agents tasked with identifying and providing such hints as auxiliary information along with answers. For example, the sleeper agent could find and share other related tables—to be either joined with (as in join discovery, e.g., (sarma2012finding)), or replacing the current table as the focus of analysis, especially if the current table proves irrelevant. Or rather than the agent having to guess why they got an empty result, the sleeper agent can provide feedback reminiscent of why-not provenance (cheney2009provenance), e.g., the probe assumed that states were encoded with two letter acronyms like “CA”, but instead they are listed out in entirety.
@@ 5.1. Supporting Exploration
Our agentic probes will consist of exploratory queries to establish grounding. Some explorations will inevitably be cast in natural language (NL) as agents may lack knowledge about the underlying databases (e.g., “how to find out how many tables are stored?”) with others expressed using SQL (e.g., SELECT count(*) FROM information_schema.tables in PostgreSQL). Today’s databases are not designed to answer NL queries. The probe optimizer in our agent-first data system must therefore orchestrate the mix of NL and SQL queries by utilizing different agents at the scale of probes.
To illustrate, consider identifying the stores that show an increasing sales trend. Our agents will first need to find out which tables are used to store sales data. A straw person probe execution plan is to pose NL questions to a web search agent to discover how to look up table schema for our specific database dialect, and execute the found queries on our database. While these are simple queries on our database’s metadata tables, the outputs returned from such queries often contain lots of unnecessary information. For instance, PostgreSQL maintains hundreds of internal tables and indexes even without any user table defined. Coupled with the user tables, the results can easily grow to thousands—or hundreds of thousands—of rows. Feeding all the rows to our query formulation agent is a waste of its limited context length. As mentioned in Section 4.1, we further need the ability to query tokens regardless of where they appear across databases, be it as part of metadata or data.
Subsequently, to discover what constitutes an increasing trend, one strategy is to find examples of “trend queries” (possibly using window queries) using NL with a web search agent, then feed the returned information to a query formulation agent to translate into SQL. We will likely get lots of example queries online, and our database will be bombarded with lots of inapplicable queries (e.g., they refer to non-existent tables, or identify the wrong trend). Worse yet, all such explorations will be mixed with other agents formulating solutions. With today’s data systems, we have no means to identify which queries are part of agentic exploration (and hence do not need to be evaluated completely). We envision that our probe optimizer will prioritize queries based on their phases (i.e., a form of admission control). Furthermore, we will store previously gleaned information using our agentic memory store to avoid repeated querying of the same information, and train agents to query our memory store instead of including such information as part of the prompt each time.
@@ 5.2.1. Intra-Probe Optimization
Deciding What to Execute. The database must first decide what queries to run and to what degree of approximation, taking the probe and its briefs into account. This requires the database to reason about the data and probe semantics, including the agent’s goals and phases. To do so, the database can use semantic query and data understanding to check if they match user’s intents, and prune away queries it deems not semantically meaningful. For instance, during the exploration phase, the database can examine the projected columns in probe to see if they are relevant to the user’s intent, and if not prune such columns, or the probes away as a whole. Moreover, the database can compare probes within a batch, guided by probe’s briefs that may have specified approximation needs across probes. The database can then make cost estimates and compare information gain from the probes to decide which probes are more helpful and/or cheaper. For example, given two probes $P$ and $P^{\prime}$ the database can prune $P^{\prime}$ away if rows that would be returned by $P^{\prime}-P$ are deemed irrelevant to the agent’s goal. This is reminiscent of prior work on pruning queries as part of visualization recommendation (seedb), and deciding query equivalence as part of query synthesis given user provided input/output examples (zloof; scythe1), although the scale of queries to compute the differences will be much larger in agentic workloads. Finally, the database can take the agent’s phase into account; for example, return coarse grain approximations during exploration, but more accurate answers during solution formulation. Beyond pruning queries, we envision agents will be able to examine other internal database states (e.g., buffer pool, outputs of query operators) to determine if it should continue with query evaluation, or move on to the next turn.
@@
Efficient Execution. As mentioned in Sec. 2, probes have substantial redundancy that we can exploit by sharing computation across them. Multi-query optimization (sharedb; mqo1; mqo2), approximate query processing (aqp) and caching of partial query results can be used to improve efficiency. However, there are new unique challenges. For example, different probes will have different approximation requirements and may be accompanied with termination criteria (a function that can be evaluated on partial results to know if they are sufficient, see Sec. 4), which makes it more difficult to reason about their semantics and what can be shared. Moreover, the database can incrementally evaluate queries, reminiscent of incremental query processing (trill), but with the new challenge of decision making across them; e.g., the database must decide which probe is the most useful to the agent and provide higher accuracy for that probe first before increasing accuracy for other probes. Finally, query planning and processing can be done jointly with optimization, e.g., the database can re-evaluate its decisions on what queries to run, or increase its level of approximation for some query during planning or processing as it obtains more information.
@@ 5.2.2. Inter-Probe Optimizations
Deciding What to Execute. Besides the strategies discussed in Section 5.2.1, the database can consider all interactions with the agent to decide what queries to run. First, it can decide on queries to run based on whether they provide any new information given past queries answered. For example, when given probes $P$ and $P^{\prime}$ by the agents across consecutive turns, if the output between $P$ and $P^{\prime}$ is expected to not provide new information to the agent—e.g., $P^{\prime}$ adds new columns that are non-relevant to the agent’s goal—then $P^{\prime}$ can be dropped. Furthermore, the database can decide what queries to run in order to minimize the number of future follow-up probes. For example, based on the agent’s goal specified in the probe briefs, it can run a query it finds maximally useful to the agent exactly and to completion rather than approximately even if the current query may take longer, expecting that the extra computation upfront will reduce total runtime across future interactions with the agent. Yet another direction is to treat the problem as one of exploration vs. exploitation: instead of always trying to provide rapid answers to queries by satisficing, the database can sometimes prioritize exploration of underexplored solution spaces to identify those solutions that have an unanticipated benefit, in order to maximize utility over time.
@@ 6. Indexing, Storage, and Transactions
On the independence front, data systems treat queries as unrelated, such that concurrent access (specifically writes) from these queries must be isolated from each other. While this simplifies application logic and ensures consistency, these mechanisms prevent cooperative sharing of state with rare exceptions (entangled). Instead of isolation, agentic workloads demand a more cooperative model—one that can safely share intermediate state across different probes, many of which are likely to be similar.
@@ 6.1. Agentic Memory Store
Updates to the Store. A separate concern is how this memory store is maintained during updates. Updates could be in the form of new probes being executed that may provide new information that augments or supersedes existing ones. Or, it could be to the underlying data or metadata, necessitating updates to any related information in the agentic memory. For example, if there is a schema update, the results of a prior probe that used that table may no longer be relevant. One approach is to allow this memory store be inconsistent with the data/metadata, and instead be updated by any new probes that discover that the information is stale. However, the downside is that the stale information may lead a new probe to make a mistake. For example, suppose the agentic memory indicates that the only relevant sales information can be found in three tables, but after that, additional relevant tables were added; here, new probes may end up returning incorrect results. Additional challenges emerge in supporting access control for multiple users. For example, agents acting on different users’ behalf could ask similar questions (e.g., ”Where is the employee’s availability stored?”). Sharing answers across such agents boosts efficiency—but raises privacy concerns, especially in the aggregate (cryptdb). Addressing these challenges will need to draw inspiration from work on knowledge bases as well as schema evolution.
@@ 6.2. Performing Branched Updates
Branch Isolation. Existing models of branching consistency, developed in the weak consistency era, e.g., in Bayou, Dynamo, or Tardis (bayou; dynamo; tardis), as well as versioned databases (huang2017orpheusdb) can offer inspiration. However, agentic speculation goes further: multiple agents may create forks that must eventually reconcile—not just with the mainline, but with each other. This requires new models of multi-agent, multi-version isolation. Most branches will be similar—e.g., same schema, 90% identical data—but isolation requires that their effects remain logically separate.
Efficient forking and rollbacks. Naively duplicating entire databases per branch is prohibitively expensive and inefficient, making support for efficient forking crucial. Industrial systems like Neon (neon), Aurora (aurora), and Bauplan (tagliabue2025safe) and academic systems like Tardis (tardis) adopt copy-on-write approaches to lazily clone state. However, these are still far from what is needed for agentic speculation at a massive scale. We need new concurrency mechanisms that exploit similarity across branches and preserve logical isolation (no cross-contamination), to enable massive parallel forking. This is analogous to MVCC on steroids: forking possibly thousands of near-identical snapshots and rolling back all but one. Unlike traditional data systems, where rollbacks are rare, we require ultra-fast rollbacks (i.e., fast aborts for failed branches).
@@ 7. Conclusion
We described our vision for data systems that natively support emerging agentic workloads. These workloads involve agentic speculation, characterized by a high-throughput, heterogeneous, and redundant mix of discovery and validation, specified by probes ideally involving a combination of queries and natural language. We present one such architecture for such redesigned data systems, and discuss emergent research challenges.
Acknowledgments
This work is supported by NSF grants IIS-1955488, IIS-2027575, DGE-2243822, IIS-2129008, IIS-1940759, and IIS-1940757, DOE awards DE-SC0016260, AC02-05CH11231, DARPA agreement HR00112590131, and funds from the state of California. This work is also supported by EPIC Data Lab sponsors and affiliates, including Adobe, Bridgewater, Google, G-Research, Microsoft, PromptQL, Sigma Computing, and Snowflake, as well as Berkeley Sky Lab sponsors and affiliates, including Accenture, AMD, Anyscale, Broadcom, Google, IBM, Intel, Intesa Sanpaolo, Lambda, Lightspeed, Mibura, Samsung SDS, SAP, Cisco, Microsoft and NVIDIA. Compute credits were provided by Azure, Modal, NSF (via NAIRR), and OpenAI. We thank the reviewers for their feedback, as well as Arash Nourian for helpful discussions.